• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 07:59:45 on October26,2001

    Entry number 75061

    keyword=owl shift summary

    At start of shift, we finished the HARP scans. Then we did a beam
    position scan in search of the posts. Concluded we are not hitting
    anything. Next, we considered doing BCM calibration but met resistance
    and decided to let day shift ask for this. Then did the target pump speed calibration and optimization. TACH=1200 chosen. Started current scan of target.
    Beam rather unstable this night.